Calculate the average of kinetic energy, in joules of the molecules in 8.0 g of methane at 27oC.
can you explain why multiply by 0.5 again in the last step?

Grade:12th pass

1 Answers

Arun
6 years ago
Total kinetic energy = n (3/2 RT0

where n = Number of moles of the gas

R = Gas constant

T = Absolute temperature

Molecular weight of methane,

CH4 = 12 + 4 * 1 = 16

∴ Number of moles of methane in 8.0 gm of methane

= 8.0/16.0 = 0.5

R = 8.314 joules/K/mole

T = 27 + 273 = 300 K

∴ Total kinetic energy of the molecules in 8.0 gm of methane at 27oC = InI * 3/2 RT

= 0.5 * 3/2 * 8.314 * 300 = 1870.65 joules

∴ Average kinetic energy = 1870.65/6.023 *1023 *0.5

= 6.21 * 10-21 joules/molecule

NOTE-  0.5 is multiplied because average kinetic  energy is asked for 8 gm methane. (molecular wt. CH4 = 16)
if for 16 gm methane, avg KE is asked then 0.5 will not be multiplied
 
h0pe it helps